OpenStudy (anonymous):

What is the opposite of 5 1/2. Sorry to ask but i really don't understand this. A.-5 1/2 B. - 2/11 C.2/11 d.2 1/5

OpenStudy (anonymous):

I think A but i am probably wrong.

OpenStudy (anonymous):

To do this, I would make it a mixed fraction. 5*2=10+1= 11.. 11/2, the opposite of 11/2 is -2/11 B

OpenStudy (anonymous):

Thank you for taking the time to explain this.

OpenStudy (anonymous):

Anytime. Message me if you ever need help again on something like this.

OpenStudy (anonymous):

Alright